# AdaCtrl: Towards Adaptive and Controllable Reasoning via Difficulty-Aware Budgeting

### v1

@@ Abstract
Figure 1: (Left) With enhanced performance (see Table 1), AdaCtrl adaptively adjusts its reasoning length, resulting in better efficiency compared to baselines; (Right) The controllable easy and hard modes of AdaCtrl allow accurate user control of reasoning budgets, as detailed in Table 2.
@@ 1 Introduction
With the emergence of test-time scaling (Snell et al., 2024; Muennighoff et al., 2025; Balachandran et al., 2025; Zhang et al., 2025), large reasoning models such as Deepseek R1 (DeepSeek-AI et al., 2025) and OpenAI O1 (OpenAI, 2024) have demonstrated superior performance across a variety of tasks by adopting more elaborate thinking strategies prior to generating final answers. While extended reasoning processes can significantly enhance a model’s ability to solve complex problems, they also introduce substantial inference overhead and may lead to the overthinking problem (Chen et al., 2025; Sui et al., 2025). For instance, even when presented with a simple problem like “Evaluate $log_{2}(64)$”, these models still tend to engage in lengthy chain-of-thought, exploring multiple solution strategies and employing advanced System-2 thinking strategies (Ryan et al., 2016; Li et al., 2025) such as planning, reflection, and verification, as the response demonstrated in right part of Figure 2.
Although such thorough reasoning is beneficial for complex queries, it results in excessive latency and resource consumption for simpler tasks, negatively impacting user experience.
To address this, recent researches have explored ways to improve reasoning efficiency. Some studies (Nayab et al., 2025; Muennighoff et al., 2025; Xu et al., 2025) leverage prompting techniques that require models to generate concise responses via explicit length constraints. Others (Munkhbat et al., 2025; Ma et al., 2025b; Xia et al., 2025b) fine-tune models on shorter reasoning trajectories. Additionally, with the growing application of reinforcement learning (RL) in LLM training, several approaches (Arora and Zanette, 2025; Aggarwal and Welleck, 2025; Hou et al., 2025) have emerged that use RL with length reward design to penalize overly long responses, thereby optimizing the model to generate outputs that are both concise and accurate. However, most existing work primarily focuses on minimizing the length of reasoning, overlooking the need for adaptive budget mechanisms that dynamically determine reasoning depth to achieve a more cost-effective balance between performance and computational efficiency. Furthermore, from the user’s perspective, a notable advantage is the controllability of the reasoning budget, which allows users to explicitly manage computational resources to prioritize either efficiency or effectiveness as needed. Despite its potential, this aspect remains relatively under-explored in current research.
To establish more adaptive and controllable reasoning approach, we propose AdaCtrl, a novel framework designed to support both adaptive and controllable reasoning. As shown in Figure 2, AdaCtrl allows the model to autonomously balance inference-time computation based on its self-assessed difficulty of the problem (i.e., adaptive reasoning mode), while also enabling users to influence the reasoning budget through designed explicit length-trigger tags (i.e. the specified easy or hard reasoning modes). Specifically, we aim to equip the model with two key meta-capabilities: (1) Self-difficulty awareness, which allows the model to estimate the difficulty of a given problem relative to its own capabilities; (2) Difficulty-aware reasoning budget allocation, which adjusts the length of the model’s response based on either the model’s own assessment or user-specified difficulty levels. To facilitate user-friendly controllable reasoning manner, we first design two length-triggered tags, namely “[Easy]” and “[Hard]”, to serve as natural interface for budget control. Then we introduce a two-stage training framework to realize these capabilities. The first stage involves cold-start fine-tuning to provide the model with an initial ability to use length-trigger tags and allocate reasoning budgets based on difficulty. The second stage employs a difficulty-aware reinforcement learning framework to refine the model’s adaptive reasoning strategies and to calibrate its self-assessment of problem difficulty according to model’s evolving capabilities during online training.
@@
Figure 2: Given same problem, AdaCtrl supports three reasoning modes, the adaptive mode dynamically allocates budgets according to the problem complexity, the easy mode offers concise answers with minimal budgets, and the hard mode delivers extensive responses with larger budgets.
Experimental results across four datasets demonstrate that AdaCtrl effectively enables models to adaptively allocate reasoning budgets through self-assessment of problem difficulty. Compared to the standard SFT + rule-based RL baseline, AdaCtrl achieves comparable accuracy on AIME2024 and yields performance improvements of 1.67% on AIME2025, 7.20% on MATH500, and 2.05% on GSM8K. Notably, these gains are accompanied by substantial reductions in response length by 10.06%, 12.14%, 62.05%, and 91.04% on the respective datasets. Moreover, AdaCtrl offers enhanced controllability via explicit length-trigger tags. When instructed to operate in the easy reasoning mode, the model reduces response lengths by 90.22% and 94.31% (from 16k to 1k tokens) on the AIME2025 and AIME2024 datasets, respectively. Conversely, under the hard reasoning mode, response lengths increase by 86.51% on GSM8K and 489.15% on MATH500. These results underscore AdaCtrl ’s capability for dynamic and interpretable control over reasoning depth. Further analysis reveals that AdaCtrl serves as an effective difficulty estimator, and accurately enables difficulty-aware budget allocation. The main contributions of this work are summarized as follows:
  • We propose AdaCtrl, which by facilitating dynamic trade-offs between efficiency and effectiveness, allows the model to emulate human cognitive behavior by modulating reasoning effort in response to varying problem demands and enhances user control on computational resources.
@@
  • We introduce a two-stage training paradigm that integrates cold-start fine-tuning and difficulty-aware reinforcement learning. This approach fosters self-awareness of task difficulty and supports difficulty-sensitive allocation of computational resources.
  • Empirical results on four benchmark datasets demonstrate that AdaCtrl successfully balances efficiency and effectiveness by adjusting to its self-assessed task difficulty. Furthermore, it offers a flexible interface that allows users to explicitly control the reasoning budget as needed.
@@ 2.1 Reasoning Efficiency via Supervised Fine-Tuning
While LLMs achieve impressive performance on complex tasks by generating elaborate multi-step reasoning chains (Dubey et al., 2024; Su et al., 2025), this capability can lead to excessive verbosity and computational overhead for simpler queries. This “overthinking” phenomenon has motivated research into improving reasoning efficiency (Qu et al., 2025a; Wang et al., 2025). One prominent strategy involves Supervised Fine-Tuning (SFT) to guide models towards more concise reasoning. Some SFT works focus on training with inherently shorter reasoning paths. For example models learn adherence to token budgets through specific prompting during data generation (Han et al., 2024). Others distill concise paths from best-of-N sampling (Munkhbat et al., 2025) or fine-tune models to omit intermediate steps for samples where the model is already confident (Yu et al., 2024). Another SFT direction compresses existing reasoning chains. Kang et al. (2024) employ GPT-4 (Achiam et al., 2023) as a compressor then fine-tune a model on these long-to-short CoT mappings. LMskip (Liu et al., 2024) induces step-skipping behavior under step constraints. SPIRIT-FT (Cui et al., 2025) identifies critical reasoning steps using perplexity as a guide for pruning. TokenSkip (Xia et al., 2025a) analyzes token importance within CoT outputs for controllable compression. These SFT methods reduce length but often enforce a general conciseness ill-suited for complex problems and typically lack self-assessment of difficulty or user budget control. While CoT-Valve (Ma et al., 2025a) offers some length modulation, our approach uniquely integrates explicit training for self-difficulty awareness and user-facing control tags within a two-stage SFT-RL framework.
2.2 Reasoning Efficiency via Reinforcement Learning
@@
RL offers another significant avenue for optimizing reasoning efficiency, building upon its success in developing deep reasoning capabilities in models like DeepSeek-Coder (Guo et al., 2025). Many RL approaches incorporate explicit length-based rewards to encourage conciseness alongside accuracy. Some methods link generation length to task difficulty or directives within the prompt: DAST (Shen et al., 2025) adapts CoT length to problem complexity via reward shaping, while LCPO (Aggarwal and Welleck, 2025) controls length using prompt-specified targets. Other techniques normalize length rewards against baselines, as seen in O1-Pruner (Luo et al., 2025), the per-prompt normalization by Arora and Zanette (2025), and the Kimi 1.5 report (Team et al., 2025). Yeo et al. (2025) proposed a cosine reward to manage length effectively, also highlighting the “length hacking” problem where models artificially extend reasoning. Beyond explicit length rewards, alternative RL strategies include meta-RL for test-time optimization (Qu et al., 2025b), utility maximization for budget awareness (Yu et al., 2025b), preference optimization with heuristics (Chen et al., 2025), and mitigating GRPO’s bias towards longer trajectories (Liu et al., 2025). However, these methods generally lack the explicit user control over reasoning depth offered by AdaCtrl’s length-trigger tags and do not prioritize training for self-awareness of problem difficulty. Our two-stage SFT-RL framework uniquely addresses these aspects, enabling both autonomous and user-influenced reasoning budgets.
@@ 3 Method
In this section, we demonstrate the design of our proposed AdaCtrl, which includes: (1) Cold-start fine-tuning that provides effective initialization for output format and difficulty estimation; (2) Difficulty-aware reinforcement learning framework that boosts model’s behavior on response length control and difficulty estimation. The whole framework is demonstrated in Figure 3.
3.1 Length-Trigger Tags as Controlling Interface
As illustrated in Figure 3, we introduce two special length-trigger tags, “[Easy]” and “[Hard]”, to more effectively utilize problem difficulty as a signal for controlling response length. Each model-generated response begins with one of these tags, indicating the anticipated complexity of the required reasoning. When presented with a problem, the model is instructed to generate a response prefaced by one of the length-trigger tags, thereby conditioning its subsequent reasoning process on the chosen tag. By equipping the model with the ability to allocate its reasoning resources according to the specified tag, we provide a straightforward interface for controlling response length: users simply supply the preferred length-trigger tag as an initial input token.
@@ 3.2 Cold-Start Fine-Tuning
This stage primarily focuses on equipping models with the ability to adhere to output formats that include length-triggered tags and to control response length accordingly. To curate suitable training data for this purpose, we directly select both easy and hard problems from the DeepMATH dataset (He et al., 2025), which provides difficulty annotations for each problem. For easy problems $\{q_{1}^{e},q_{2}^{e},...,q_{n}^{e},\}$, we utilize the model to be trained $\mathcal{M}$ (i.e. Qwen2.5-7B-Instruct) to generate concise response, while for hard ones $\{q_{1}^{h},q_{2}^{h},...,q_{m}^{h},\}$, a strong large reasoning model $\mathcal{R}$ is employed to generate reasoning trace. Then we filter the correct part for both subset:
$\mathcal{D}^{e}=\{(q_{i}^{e},[r_{i},y_{i}])|[r_{i},y_{i}]=\mathcal{M}(q_{i}^{e}),\mathbb{I}(y_{i},\hat{y}_{i})=1\},$ (1)
$\mathcal{D}^{h}=\{(q_{i}^{h},[r_{i},y_{i}])|[r_{i},y_{i}]=\mathcal{R}(q_{i}^{h}),\mathbb{I}(y_{i},\hat{y}_{i})=1\},$ (2)
where $[r_{i},y_{i}]$ is the model-generated response that includes reasoning process $r_{i}$ and the predicted answer $y_{i}$, $\hat{y}_{i}$ is the ground truth answer of corresponding samples, and $\mathbb{I}(y_{i},\hat{y}_{i})=1$ represents that the model response $[r_{i},y_{i}]$ is correct with the verification of answer $\hat{y}_{i}$.
Figure 3: AdaCtrl comprises a two-stage training pipeline: the cold-start finetuning first utilizes both short and long reasoning trajectories to establish basic budget awareness, then a difficulty-aware reinforcement learning framework is utilized to calibrate problem difficulty estimation and develop adaptive reasoning strategies.
By prepending the “[Easy]” tag for response in $\mathcal{D}^{e}$ and “[Hard]” for those in $\mathcal{D}^{h}$, we construct a dataset $\mathcal{D}=\{(q_{i},[t_{i},r_{i},y_{i}])\}$ for supervised fine-tuning (SFT) as cold start, where $t_{i}$ is length-trigger tag. After the cold start finetuning, model generated responses will start with above length-trigger tags.
@@ 3.3 Difficulty-Aware Reinforcement Learning
We utilize GRPO (Shao et al., 2024) as the training algorithm, which produce multiple rollouts that can be naturally utilized to estimate difficulty of given question from the perspective of the trained model.
3.3.1 Reward Design
Outcome Accuracy Reward. The rule-based outcome accuracy reward has been widely utilize in RL training Guo et al. (2025); Aggarwal and Welleck (2025); Arora and Zanette (2025); Yu et al. (2025a), which evaluates the correctness of a generated response:
@@ 3.3.1 Reward Design
Difficulty Estimation Calibration Reward. During reinforcement learning (RL) training, the model’s capability evolves dynamically. Same questions may need different length-trigger tags at different training steps. To account for this, we calibrate the model’s estimated difficulty using the accuracy of multiple rollouts. Through a pre-defined threshold $\delta$, we label a question as easy if the rollout accuracy exceeds $\delta$, or it is labeled as hard. During rollout process, the model is expected to generate a correct length-trigger tag at beginning of whole response as correct difficulty estimation, so the reward function is designed based on the matches between length-trigger tag $t_{i}$ and estimated tag label $\hat{t}_{i}$ that determined by multiple rollouts:
@@
Difficulty-aware Length Reward. Different from previous related works that encourage the model to generate concise responses for all problems, we hope to encourage such behavior only for easy problems and maintain long thinking capabilities for better tackling hard problems. To prevent unnecessary overthinking, we design a difficulty-aware length reward that encourage concise responses when the generated length-trigger tag $t_{i}$ is “[Easy]”:
@@
where $l_{i}^{j}$ is the $j$-th rollout length for problem $q_{i}$, and $L_{i}$ is the max length in the rollout group of problem $q_{i}$. Utilizing the monotonicity of cosine function within a specific domain (i.e. 0 - $\pi$), the difficulty-aware length reward will assign lower score with the increase of generated response length.
3.3.2 Reinforcement Learning for Difficulty-Aware Budgeting
To enhance the model’s adaptive reasoning capabilities and calibrate its self-evaluation of problem difficulty during online training, the overall reward in the reinforcement learning process is computed by integrating three specifically designed reward components:
@@ 3.3.2 Reinforcement Learning for Difficulty-Aware Budgeting
where $\alpha$ and $\beta$ are hyper-parameters.
At this stage, we sampled another collection of problems that covers both easy and hard problems from DeepMATH (He et al., 2025) to allow the model to learn dynamic strategies for different types of problems. Following Yu et al. (2025a); He et al. (2025), the policy model $\pi_{\theta}$ is optimized through the following objective:
@@
$\ \text{clip}\Big{(}\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid q,o_{i,<t})}(\theta),1-\varepsilon,1+\varepsilon\Big{)}\hat{A}_{i,t}\Big{)}-\beta D_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})\Bigg{)}\Bigg{]},$
where $G$ is the group size, $o_{i}$ is the rollout response, $\hat{A}_{i,t}$ is the advantage of the $i$-th response calculated by normalizing rewards in the group, and $\beta,\varepsilon$ are hyper-parameters.
Through reinforcement learning, models can more effectively assess problem difficulty relative to their own capabilities and develop adaptive reasoning strategies accordingly.

### v2

@@ 1 Introduction
With the emergence of test-time scaling (Snell et al., 2024; Muennighoff et al., 2025; Balachandran et al., 2025; Zhang et al., 2025), large reasoning models such as Deepseek R1 (DeepSeek-AI et al., 2025) and OpenAI o1 (OpenAI, 2024) have demonstrated superior performance across a variety of tasks by thoroughly exploring various reasoning paths prior to generating final answers. While such deep and long reasoning processes can significantly enhance a model’s ability to solve complex problems, they inevitably introduce significant inference overhead and lead to unnecessary overthinking on simpler problems (Chen et al., 2025; Sui et al., 2025). For instance, even when presented with easy and straightforward questions like “Evaluate $log_{2}(64)$”, these models still tend to engage in lengthy chain-of-thought (Wei et al., 2022), unnecessarily employing advanced meta-reasoning skills such as planning, reflection, and verification (Ryan et al., 2016; Li et al., 2025; Wang et al., 2025a). Such reasoning behavior, while beneficial for complex queries, incurs excessive latency and computational costs, negatively affecting user experience (Wang et al., 2025c).
@@
Figure 1: Given the same problem, AdaCtrl supports three reasoning modes: the easy mode offers concise answers with less tokens; the hard mode delivers extensive responses with more tokens; and the adaptive mode dynamically allocates reasoning budgets according to the problem complexity.
Recent efforts have explored several ways to improve the reasoning efficiency and mitigate overthinking issue. Some approaches aim to minimize reasoning length across all questions, regardless of their actual complexity (e.g., easy or hard), by enforcing conciseness through fine-tuning (Munkhbat et al., 2025; Ma et al., 2025) or reinforcement learning (Arora and Zanette, 2025; Aggarwal and Welleck, 2025). However, this universal compression strategy may sacrifice reasoning quality on truly complex tasks that require deeper analysis. On the other hand, some methods attempt to precisely control the token budget for each question, seeking a more efficient yet accurate reasoning process (Nayab et al., 2025; Muennighoff et al., 2025; Xu et al., 2025). Nevertheless, such fine-grained control often relies on accurate difficulty assessments and exceptional capabilities of underlying LLMs (e.g., precise instruction-following), which can be inflexible and brittle in practice. It is inherently challenging to determine the optimal token budget beforehand, and models may struggle to reliably interpret or execute fine-grained instructions. These limitations underscore the need for a unified framework that is both adaptive and easily controllable based on task complexity.
To this end, we propose AdaCtrl, a novel framework to 1) allow the model to dynamically adjust their reasoning effort based on self-assessed difficulty (i.e., adaptive reasoning mode); and 2) enable users to specify reasoning mode according to difficulty of problem (i.e., controllable budget allocation). As shown in Figure 1, given a query, AdaCtrl offers three reasoning modes. Two of these modes are manually specified to allow users to control the reasoning budget: (i) Easy, which prioritizes efficiency and provides concise responses to any question; (ii) Hard, which aims for higher effectiveness by elaborating the full reasoning process and delivering detailed information. In addition, an adaptive mode automatically adjusts the reasoning effort based on the complexity of the input query, achieving a balance between effectiveness and efficiency without manual intervention. To achieve this, we start by curating a dataset that covers both easy and hard subsets and insert corresponding indicator tags (i.e., “[Easy]” or “[Hard]”) prior to the model’s responses. Then we utilize cold-start fine-tuning to empower the model with foundation capabilities to estimate the complexity of a question and allocate reasoning budgets accordingly, rendering self-aware difficulty estimation. Moreover, we adapt difficulty-aware reinforcement learning framework with carefully designed rewards to calibrate its self-assessment of problem difficulty (i.e., difficulty estimation calibration reward) and to refine the model’s adaptive reasoning behavior (i.e., difficulty-aware length reward).
Experimental results on four benchmarks demonstrate that AdaCtrl significantly improves the trade-off between effectiveness and efficiency. It outperforms most baselines across four datasets while efficiently managing the reasoning budget. Specifically, compared with the standard SFT + RL baselines, AdaCtrl achieves accuracy improvements of up to 10.14%, while reducing response length by as much as 91.04%. Further analysis demonstrates that AdaCtrl offers effective human-in-the-loop controllability via explicit difficulty-aware tags, enables accurate difficulty estimation during reinforcement learning, and maintains robust performance under hyperparameter variation. Overall, the contributions are as follows:
  • We introduce AdaCtrl, a unified framework for adaptive and controllable reasoning that supports dynamic trade-offs between efficiency and performance, allowing the model to estimate the difficulty of problem and adjust the reasoning mode itself, and also the user to specify the desired reasoning mode to meet diverse needs in practice.
@@
  • We present a two-stage training paradigm that integrates cold-start fine-tuning and difficulty-aware reinforcement learning together to foster self-awareness of problem difficulty and supports difficulty-aware budget allocation, considering the differences and dynamics of model capabilities.
  • Empirical results on four benchmark datasets demonstrate that AdaCtrl successfully enhance adaptivity and controllability via explicit difficulty-aware tags. Further analysis reveals that AdaCtrl serves as an effective difficulty estimator, and accurately enables difficulty-aware budget allocation.
@@ 2 Related Work
Reasoning Efficiency via Supervised Fine-Tuning. While LLMs achieve impressive performance on complex tasks by generating elaborate multi-step reasoning chains (Dubey et al., 2024; Su et al., 2025), this capability can lead to excessive verbosity and computational overhead for simpler queries. This “overthinking” phenomenon has motivated research into improving reasoning efficiency (Qu et al., 2025a; Wang et al., 2025c). One prominent strategy involves Supervised Fine-Tuning (SFT) to guide models towards more concise reasoning. Some SFT works focus on training with inherently shorter reasoning paths. For example models learn adherence to token budgets through specific prompting during data generation (Han et al., 2024). Others distill concise paths from best-of-N sampling (Munkhbat et al., 2025) or fine-tune models to omit intermediate steps for samples where the model is already confident (Yu et al., 2024). Another SFT direction compresses existing reasoning chains. Kang et al. (2024) employ GPT-4 (Achiam et al., 2023) as a compressor then fine-tune a model on these long-to-short CoT mappings. LMskip (Liu et al., 2024) induces step-skipping behavior under step constraints. SPIRIT-FT (Cui et al., 2025) identifies critical reasoning steps using perplexity as a guide for pruning. TokenSkip (Xia et al., 2025) analyzes token importance within CoT outputs for controllable compression. These SFT methods reduce length but often enforce a general conciseness ill-suited for complex problems and typically lack self-assessment of difficulty or user budget control.
@@
Reasoning Efficiency via Reinforcement Learning. RL offers another significant avenue for optimizing reasoning efficiency, building upon its success in developing deep reasoning capabilities in models like DeepSeek-Coder (Guo et al., 2025). Many RL approaches incorporate explicit length-based rewards to encourage conciseness alongside accuracy. Some methods link generation length to task difficulty or directives within the prompt: DAST (Shen et al., 2025) adapts CoT length to problem complexity via reward shaping, while LCPO (Aggarwal and Welleck, 2025) controls length using prompt-specified targets. Other techniques normalize length rewards against baselines, as seen in O1-Pruner (Luo et al., 2025), the per-prompt normalization by Arora and Zanette (2025), and the Kimi 1.5 report (Team, 2025). Yeo et al. (2025) proposed a cosine reward to manage length effectively, also highlighting the “length hacking” problem where models artificially extend reasoning. Beyond explicit length rewards, alternative RL strategies include meta-RL for test-time optimization (Qu et al., 2025b), utility maximization for budget awareness (Yu et al., 2025b), preference optimization with heuristics (Chen et al., 2025), and mitigating GRPO’s bias towards longer trajectories (Liu et al., 2025). However, these methods generally lack the explicit user control over reasoning depth offered by AdaCtrl’s difficulty-aware tags and do not prioritize training for self-awareness of problem difficulty. Our two-stage SFT-RL framework uniquely addresses these aspects, enabling both autonomous and user-influenced reasoning budgets.
Figure 2: AdaCtrl comprises a two-stage training pipeline: the cold-start finetuning first utilizes both short and long reasoning trajectories to establish basic budget awareness; then a difficulty-aware reinforcement learning framework is utilized to calibrate problem difficulty estimation and develop adaptive reasoning strategies.
@@ 3 Method
In this section, we demonstrate the design of our proposed AdaCtrl, which includes: (1) Cold-start fine-tuning that provides initialization for difficulty estimation and difficulty-aware budget adjustment; (2) Difficulty-aware reinforcement learning framework that boosts model’s capabilities on response length control and difficulty estimation. The entire framework is demonstrated in Figure 2.
@@ 3.1 Cold-Start Fine-Tuning
This stage primarily focuses on equipping models with the ability to adhere to output formats that include difficulty-aware tags (e.g., "[Easy]” and "[Hard]”) [^2] We do not consider more fine-grained category in order to maintain ease of control and ensure reliability. and to control response length accordingly. To curate suitable training data for this purpose, we directly select both easy and hard problems from the DeepMATH dataset (He et al., 2025), which provides difficulty annotations for each problem. For easy problems $\{q_{1}^{e},q_{2}^{e},...,q_{n}^{e},\}$, we utilize the instruction model, $\mathcal{M}$, to generate concise response, while for hard ones $\{q_{1}^{h},q_{2}^{h},...,q_{m}^{h},\}$, a strong large reasoning model $\mathcal{R}$ is employed to generate reasoning trace. Then we filter the solutions based on answer correctness for both subsets:
$\mathcal{D}^{e}=\{(q_{i}^{e},[p_{i},y_{i}])|[p_{i},y_{i}]=\mathcal{M}(q_{i}^{e}),\mathbb{I}(y_{i},\hat{y}_{i})=1\},$ (1)
$\mathcal{D}^{h}=\{(q_{i}^{h},[p_{i},y_{i}])|[p_{i},y_{i}]=\mathcal{R}(q_{i}^{h}),\mathbb{I}(y_{i},\hat{y}_{i})=1\},$ (2)
where $r_{i},y_{i}]$ is the model-generated response that includes reasoning process $p_{i}$ and the predicted answer $y_{i}$, $\hat{y}_{i}$ is the ground truth answer of corresponding samples, and $\mathbb{I}(y_{i},\hat{y}_{i})=1$ represents that the model response $p_{i},y_{i}]$ is correct with the verification of answer $\hat{y}_{i}$. Then by prepending the “[Easy]” tag for response in $\mathcal{D}^{e}$ and “[Hard]” tag for those in $\mathcal{D}^{h}$, we construct a dataset $\mathcal{D}=\{(q_{i},[t_{i},p_{i},y_{i}])\}$ for cold-start training, where $t_{i}$ is the difficulty-aware tag. After the cold-start finetuning, the model learns to adhere to the specified format and to generate solutions based on its estimated difficulty of each problem, such as allocating more reasoning tokens for hard problems to enable deeper and more diverse reasoning paths. Notably, this also provides greater controllability for non-expert users, who can easily guide the model’s behavior by using simple tags like “[Easy]” and “[Hard]”.
@@ 3.2 Difficulty-Aware Reinforcement Learning
After cold-start fine-tuning, we adapt Group Relative Policy Optimization (GRPO) (Shao et al., 2024) as the reinforcement learning (RL) algorithm because it produces multiple rollouts that can be naturally utilized to estimate difficulty of given question from the perspective of the trained model. Specifically, we carefully design three types of rewards to guide the difficulty-aware reinforcement learning optimization:
Outcome Accuracy Reward.
The rule-based outcome accuracy reward has been widely utilize in RL training (Guo et al., 2025; Aggarwal and Welleck, 2025; Arora and Zanette, 2025; Yu et al., 2025a; Wang et al., 2025b), which evaluates the correctness of a generated response:
@@
Difficulty Estimation Calibration Reward.
Accurately estimating problem difficulty is a fundamental aspect of difficulty-aware budgeting. However, since the difficulty of a problem can vary significantly across different models, and obtaining reliable, model-specific difficulty labels is also challenging. To address this, we design a reward function that leverages GRPO’s multiple rollouts to estimate a golden difficulty during online training in a natural and effective manner. Specifically, we calibrate the estimated difficulty of the model using the frequency of rollouts that leads to a correct answer. We label a question as easy if the frequency of accurate rollouts exceeds a pre-defined threshold $\delta$, or it is labeled as hard. Moreover, during the rollout process, the model is expected to generate a correct difficulty-aware tag at the beginning of the entire response to match the its capability, so the reward function is designed based on the matches between the generated difficulty-aware tag $t_{i}$ and estimated tag label $\hat{t}_{i}$ that is determined by multiple rollouts:
@@
Difficulty-aware Length Reward.
Different from previous related works that encourage the model to generate concise responses for all problems (Arora and Zanette, 2025; Aggarwal and Welleck, 2025), we hope to encourage such behavior only for easy problems and maintain long thinking capabilities for better tackling hard problems. This helps prevent over-optimization, where the model may fail to engage in deeper thinking when necessary. To prevent unnecessary overthinking, we design a difficulty-aware length reward that encourage concise responses when the generated difficulty-aware tag $t_{i}$ is “[Easy]”:
@@
[…]
Overall Reward and Objective.
To enhance the model’s adaptive reasoning capabilities and calibrate its self-evaluation of problem difficulty during online training, the overall reward in the reinforcement learning process is computed by integrating the three specifically designed rewards:
@@
where $\alpha$ and $\beta$ are hyper-parameters. During optimization, we sample a collection of problems that covers both easy and hard problems from DeepMATH (He et al., 2025) to allow the model to learn dynamic strategies for different types of problems. Following Yu et al. (2025a); He et al. (2025), the policy model $\pi_{\theta}$ is optimized through the following objective:
@@
$\ \text{clip}\Big(\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid q,o_{i,<t})}(\theta),1-\varepsilon,1+\varepsilon\Big)\hat{A}_{i,t}\Big)-\beta^{\prime}D_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})\Bigg)\Bigg],$
where $G$ is the group size, $o_{i}$ is the rollout response, $\hat{A}_{i,t}$ is the advantage of the $i$-th response calculated by normalizing rewards in the group, and $\beta^{\prime},\varepsilon$ are hyper-parameters. Through this reinforcement learning process, models can more effectively assess problem difficulty relative to their own capabilities and develop adaptive reasoning strategies accordingly.